(Make a chart and solve) A chemist has two solutions, Solution A is 8% acetic acid and Solution B is 20% acetic acid. If he wants 4 liters of an 11% acetic acid solution, how much of solution A and solution B should he mix ?
Answer by Alan3354(69443)

(Make a chart and solve) A chemist has two solutions, Solution A is 8% acetic acid and Solution B is 20% acetic acid. If he wants 4 liters of an 11% acetic acid solution, how much of solution A and solution B should he mix ?
====================
A + B = 4 --- total solution
8A + 20B = 11*(A + B) --- total acid
etc.
------
IDK why a chart is needed.